A beam of vertically polarized linear light is perpendicularly incident on an ideal linear polarizer. Show that if its transmission axis makes an angle of 60° with the vertical only 25% of the irradiance will be transmitted by the polarizer. 


1
Expert's answer
2021-04-29T07:28:14-0400

Let the irradiance of beam of polarized light be "I(0)" then,

According to Malus Law,


"I(\theta)=I(0)cos^2\theta\ \ \ \ .....(i)"

where "\theta" = angle between direction of polarized light and transmission axis

Here "\theta" = "{60}\degree"

"I(\theta)=" Irradiance of polarized light after emerging from ideal linear polarizer


On substituting the value of "\theta" = "{60}\degree" in equation (i) we get,


"I(60\degree)=I(0)cos^260\degree\\I(60\degree)=I(0)\times(0.5)^2\\I(60\degree)=0.25I(0)\\\Rightarrow I(60\degree)=25\% \ of \ \ I(0)"

Hence, 25% of the irradiance will be transmitted by the polarizer if transmission axis makes an angle of 60° with the vertical .